OpenStudy (anonymous):

Reggie is three less than four times his daughter's age. Reggie is 69; how old is his daughter? 10 18 12 20

OpenStudy (anonymous):

69/4 - 3= 14.25

OpenStudy (anonymous):

?

OpenStudy (anonymous):

thatas not an option tho

OpenStudy (anonymous):

sorry. 69+3= 72. so 72/4= 18 awnser is 18

OpenStudy (anonymous):

thnks